OpenStudy (anonymous):

Need some help please...What is the radius for the circle given by the quation...(x-2)^2+y^2=14..Round your answer to the nearest thousandth.

OpenStudy (anonymous):

general equation of circle: \[(x-h)^{2} + (y-k)^{2} = r^{2}\] the center is the point (h,k) and the radius is r. r = sqrt(r^2)

OpenStudy (anonymous):

so now what should i do

OpenStudy (anonymous):

well. what is r^2 of your circle if you compare it to the general equation?

OpenStudy (anonymous):

i'm not sure i hate conic sections :(

OpenStudy (anonymous):

the equation is already in it's general form. the center is at (2, 0) and r^2 = 14. to find r you'll need a calculator

OpenStudy (anonymous):

i'm not at school i dont have one :/

OpenStudy (anonymous):

if you're on the internet you do :). or even just on a computer. i bet you have a cellphone in your pocket :P. regardless. sqrt14 = 3.741657387

OpenStudy (anonymous):

i thought i would need a scientific one :( how could i round tht to the nearest thousandth?

OpenStudy (anonymous):

the thousanth's in the third decimal (tenths, hundreths, thousandths, ten thousannths, ect).

OpenStudy (anonymous):

to round you look at the number after the thousanths. if its > 5 you round up. < 5, you leave as is

OpenStudy (anonymous):

im so lost

OpenStudy (anonymous):

the answer is 3.742. we rounded up

OpenStudy (anonymous):

sorry im such a pain lol

OpenStudy (anonymous):

lol no problem. i'm here to help

OpenStudy (anonymous):

thank you :)